# IFJPAN-IV-2023-3

### v1

@@ 5.2 The effect of $Z$-boson exchange at Belle II energy
Let us turn our attention to lower energies. In Ref. [11], we studied effects of the dipole form-factors $A(s)$ and $B(s)$ on spin effects and on an observable of potential interest for the Belle II measurement. Since this is expected to be a precision measurement, that is why the effect of $Z$-boson exchange, or better to say the $Z-\gamma^{*}$ interference, which could represent potential bias, should be addressed. In Fig. 5, we present the ratio of distributions where the $Z$-boson contribution is turned on or off. The distributions are prepared exactly in the same manner as for Fig. 2 of Ref. [11]. The $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ decay channels are used for the production process $e^{-}e^{+}\to\tau^{-}\tau^{+}n\gamma$ at Belle II energy. Acoplanarity between the planes spanned by the $\pi^{-}\pi^{0}$ system (from $\tau^{-}$) and the $\pi^{+}\pi^{0}$ system (from $\tau^{+}$), respectively, is monitored. The momenta are boosted to the rest frame of visible decay product system of $\pi^{-}\pi^{0}\pi^{+}\pi^{0}$ mesons. An additional selection criteria on the pions energies, $(E_{\pi^{-}}-E_{\pi^{0}})(E_{\pi^{+}}-E_{\pi^{0}})$ to be larger or smaller than zero, is applied.
@@
Figure 5: Impact of $Z$-exchange to the acoplanarity of $\pi^{\pm}\pi^{0}$ planes of $\tau^{\pm}$ decay products. The production process $e^{+}e^{-}\to\tau^{+}\tau^{-}n\gamma$, and decays $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ at the center-of-mass energy $10.58$ GeV are used. Events are selected so that there is the same sign of energy differences for the charged and neutral pions coming from decays $\tau^{-}\to\pi^{-}\pi^{0}\nu_{\tau}$ and $\tau^{+}\to\pi^{+}\pi^{0}\bar{\nu}_{\tau}$. The detailed definition of presented observable is given in Ref. [11]. The ratio of the acoplanarity distributions with and without $Z$-boson contribution is shown.
@@ 6 Summary and outlook
These studies include calculation of the analytical formulas for spin-amplitude components and spin-correlations matrices, for $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes. In case of the $s$-channel $f_{i}\,\bar{f}_{i}\to\tau^{-}\tau^{+}$ process, including photon and $Z$-boson exchange and their interference makes those formulas applicable for a large range of $\tau$-pair invariant mass, from the Belle II energies to $WW$ and $ZZ$ thresholds. Numerical predictions for the ratios of transverse components of spin-correlations matrices were also shown and discussed.
@@
To demonstrate applications of the re-weighting algorithm, semi-realistic observables are suggested, and for $e^{-}e^{+}$ case numerical results are presented for anomalous dipole moments at high energies and for $Z$-boson exchange contribution to Belle II energies.
@@
The algorithm for the implementation of anomalous magnetic and electric dipole moments for the s-channel photon exchange in $e^{+}e^{-}\to\tau^{+}\tau^{-}$ process was enriched with contribution of $Z$-boson exchange. With the help of the extension to include $Z$-boson contributions, the impact on dipole moment observable ambiguities at Belle II energies can be studied. The algorithm can be used also for higher energies in its present form, up to about $ZZ$-boson pair production threshold. Above that scale, contributions from $WW$ and $ZZ$-boson pair production become double resonant and box diagram contribution became sizable. Their impact on spin effects can not be neglected. The observable, acoplanarity of decay products of $\tau$ lepton pairs, survive upgrade to high energies after minor modifications, provided the effects of photon radiation can be ignored. We have checked the results for a new observable, and also with the help of analytic calculations. We can conclude that tools for further studies of more experimental details is prepared. It is again an add-up to KKMC and may be useful not only for the FCC oriented studies, but also for the evaluation of potential biases in the Belle II precision measurements as well.
@@
Reliability of reweighting at high energies require some attention, but as we work for a discovery tool, not a precision tool, that may be postponed to further studies. So far, test when photon radiation is removed with generation variable cuts were performed only. That is suitable for Belle II applications, but for high energy applications, such as at the FCC, further work on selection cuts to prevent degradation of the performance due to bremsstrahlung is needed. It will require careful future considerations for detector acceptance and more focus on the details of the detector than on tool validation. At the FCC, even at the center-of-mass energy around the $Z$-boson peak, bremsstrahlung photons lost in the beam pipe may have $p_{T}$ comparable to the $\tau$ lepton mass. That means further studies and program extensions are needed for such configurations. However, the program can now be used for studies with no bremsstrahlung events or events of rather soft photons with energy and/or $p_{T}$ sizably smaller than half of the $\tau$ mass. This is not a constraint for program use at Belle II energies, where bremsstrahlung photons lost in the beam pipe cannot have large $p_{T}$. At high energies, the so-called Mustraal frame [36, 37, 30] for weight calculation may need to be used, as already proposed for evolution of future versions of TauSpinner.
@@ Acknowledgments
A.Yu.K. acknowledges partial support by the National Academy of Sciences of Ukraine via the program “Participation in the international projects in high-energy and nuclear physics” (project C-4/53-2023).
This project was supported in part from funds of Polish National Science Centre under decision DEC-2017/27/B/ST2/01391 and of COPIN-IN2P3 collaboration with LAPP-Annecy.
This project was supported by the U.S. Department of Energy under research Grant No. DE-SC0022350.

### v2

@@ 5.2 The effect of $Z$-boson exchange at Belle II energy
Let us turn our attention to lower energies. In Ref. [11], we studied effects of the dipole form-factors $A(s)$ and $B(s)$ on spin effects and on an observable of potential interest for the Belle II measurement. Since this is expected to be a precision measurement, that is why the effect of $Z$-boson exchange, or better to say the $Z\gamma$ interference, which could represent potential bias, should be addressed. In Fig. 5, we present the ratio of distributions where the $Z$-boson contribution is turned on or off. The distributions are prepared exactly in the same manner as for Fig. 2 of Ref. [11]. The $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ decay channels are used for the production process $e^{-}e^{+}\to\tau^{-}\tau^{+}n\gamma$ at Belle II energy. Acoplanarity between the planes spanned by the $\pi^{-}\pi^{0}$ system (from $\tau^{-}$) and the $\pi^{+}\pi^{0}$ system (from $\tau^{+}$), respectively, is monitored. The momenta are boosted to the rest frame of visible decay product system of $\pi^{-}\pi^{0}\pi^{+}\pi^{0}$ mesons. An additional selection criteria on the pions energies, $(E_{\pi^{-}}-E_{\pi^{0}})(E_{\pi^{+}}-E_{\pi^{0}})$ to be larger or smaller than zero, is applied.
@@
Figure 5: Impact of $Z$-exchange to the acoplanarity of $\pi^{\pm}\pi^{0}$ planes of $\tau^{\pm}$ decay products. The production process $e^{+}e^{-}\to\tau^{+}\tau^{-}n\gamma$, and decays $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ at the c.m. energy $10.58$ GeV are used. Events are selected so that there is the same sign of energy differences for the charged and neutral pions coming from decays $\tau^{-}\to\pi^{-}\pi^{0}\nu_{\tau}$ and $\tau^{+}\to\pi^{+}\pi^{0}\bar{\nu}_{\tau}$. The detailed definition of presented observable is given in Ref. [11]. The ratio of the acoplanarity distributions with and without $Z$-boson contribution is shown.
@@ 6 Summary and outlook
These studies include calculation of the analytical formulas for spin-amplitude components and spin-correlations matrices, for $e^{-}e^{+}\to\tau^{-}\tau^{+}$, $q\bar{q}\to\tau^{-}\tau^{+}$ and $\gamma\gamma\to\tau^{-}\tau^{+}$ processes. In case of the $s$-channel $f_{i}\,\bar{f}_{i}\to\tau^{-}\tau^{+}$ process, including photon and $Z$-boson exchange, their interference and radiative corrections in framework of IBA, makes those formulas applicable for a large range of $\tau$-pair invariant mass, from the Belle II energies up to above $WW$ and $ZZ$ pair production thresholds. Numerical predictions for the ratios of transverse components of spin-correlations matrices were also shown and discussed.
@@
To demonstrate applications of the reweighting algorithm, semi-realistic observables are suggested, and for $e^{-}e^{+}$ case numerical results are presented for anomalous dipole moments at high energies and for $Z$-boson exchange contribution to Belle II energies.
@@
The algorithm for the implementation of anomalous magnetic and electric dipole moments for the s-channel photon exchange in $e^{+}e^{-}\to\tau^{+}\tau^{-}$ process was enriched with contribution of $Z$-boson exchange. With the help of the extension to include $Z$-boson contributions, the impact on dipole moment observable ambiguities at Belle II energies can be studied. The algorithm can be used also for higher energies in its present form up to and above $ZZ$-boson pair production threshold. At these energies, contributions from $WW$- and $ZZ$-boson pair production become double resonant and box diagram contributions become sizable. Their impact on spin effects cannot be neglected.
The observable, acoplanarity of decay products of $\tau$ lepton pairs, survive upgrade to high energies after minor modifications, provided the effects of photon radiation can be ignored, as in the case of $\sqrt{s}\simeq M_{Z}^{2}$. We have checked the results for a new observable, and also with the help of analytic calculations. We can conclude that tools for further studies of more experimental details is prepared. It is again an add-up to KKMC and may be useful not only for the FCC oriented studies, but also for the evaluation of potential biases in the Belle II precision measurements as well.
@@
There is important issue about the QCD and EW corrections; those corrections have been extensively studied, for example, in Refs. [38, 39, 40, 41], in which progress in automated calculation of matrix elements has been achieved. Of course, for precise SM simulation of the $\tau$-pair production, these higher-order corrections need to be taken into account. In the present paper, these corrections are included in framework of improved Born approximation [25, 17].
Also the effects of multi-parton interactions [42, 43] and underlying events [44, 45, 46] may be of importance. Note that in general, evaluations of underlying events and multi-parton interaction effects are investigated in framework of experimental collaboration work. At the same time, experimental details of detection and background subtractions have to be taken into account. Let us stress again that all these efforts are needed to assure robustness of the event sample. Only if its precision is established, one can expect reweighting results to be meaningful.
The TauSpinner reweighting algorithm for introduction of new effects is expected to supplement events with rather small corrections: event weights close to 1. That is why demand on precision of NP effects is not very strict, provided that predictions of the SM are not compromised. The use of the present paper algorithms may be considered as the first step of the work, as we assume that the impact of interactions resulting with dipole moments will not affect the SM part of the interaction sizably.
@@
Reliability of reweighting at high energies requires some attention, but as we work for a discovery tool, not a precision tool, that may be postponed to further studies. So far, tests when photon radiation is removed with generation variable cuts were performed only. That is suitable for Belle II applications, but for high energy applications, such as at the FCC, further work on selection cuts to prevent degradation of the performance due to bremsstrahlung is needed. It will require careful future considerations for detector acceptance and more focus on the details of the detector than on tool validation. At the FCC, the bremsstrahlung photons lost in the beam pipe may have $p_{T}$ comparable to the $\tau$-lepton mass. That means further studies and program extensions are needed for such configurations. For example, at the c.m. energy around the $Z$-boson peak presence of such photons is largely reduced. However, the program can now be used for studies with no bremsstrahlung events or events of rather soft photons with energy and/or $p_{T}$ sizably smaller than half of the $\tau$ mass. This is not a constraint for program use at Belle II energies, where bremsstrahlung photons lost in the beam pipe cannot have large $p_{T}$. At high energies, the so-called Mustraal frame [47, 48, 33] for weight calculation may need to be used, as already proposed for evolution of future versions of TauSpinner.
@@ Acknowledgments
A.Yu.K. acknowledges partial support by the National Academy of Sciences of Ukraine via the program “Participation in the international projects in high-energy and nuclear physics” (project No. C-4/53-2023). He is grateful to the Polish Academy of Sciences for financial support during his stay at the Institute of Nuclear Physics PAS. This project was supported in part from funds of the Polish National Science Centre under Decisions No. DEC-2017/27/B/ST2/01391, No. UMO-2022/01/3/ ST2/00027, and of COPIN-IN2P3 collaboration with Laboratoire d’Annecy de Physique des Particules (LAPP), Annecy. This project was supported by the U.S. Department of Energy under research Grant No. DE-SC0022350.